OpenStudy (anonymous):

What is the probability of drawing a diamond from a standard deck of cards on a second draw, given that a diamond was drawn on the first draw and not replaced? twelve over fifty one one over fifty two one over fifty one thirteen over fifty two

OpenStudy (anonymous):

the ratio of the remaining number of diamonds to the remaining number of cards

OpenStudy (anonymous):

12/51

OpenStudy (anonymous):

In a normal deck of cards there are 52 cards, 13 of which are diamonds. (52/4=13). BUT we already drew a diamond and didn't put it back so there are noe 51 cards, 12 of which are diamonds. That means we have 51 total possibilities, 12 of which are correct. To find the probability we do \[\frac{correct}{total}=\frac{12}{51}\] which is also about 0.235 or 23.5%

OpenStudy (anonymous):

ty can you help me with this one A diner asked its customers, "Did you cook dinner last night?" The results on the survey are shown in the table below: Male Female Cooked dinner 312 234 Did not cook dinner 146 454 What is the probability that one of the customers chosen from this survey was a female and did not cook dinner?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

0.51 0.65 0.76 0.40

OpenStudy (anonymous):

Same idea as before. We already know the number of "correct" answers (females who didn't cook dinner) is 454. We also need to know the total number of people. We can get this by adding up all the numbers. 312+234+146+454=1146. So again we can do: \[\frac{correct}{total}=\frac{454}{1146}=0.39616\] which is 0.40 when rounded to two decimal palces.

OpenStudy (anonymous):

Male Female Cooked dinner 312 234 Did not cook dinner 146 \(\color{red}{454}\)

OpenStudy (anonymous):

ty so much @twvogels